Need to answer this question using energy equations (physics 12)

Question

What must be the initial velocity if point y was 13.5m high and 500J of energy was lost to friction? The cart and people have a mass of 400 kg and the final velocity is still 10 m/s.


1
Expert's answer
2021-03-29T08:56:55-0400
"0.5m(v_i^2-v_f^2)+mgh=W_f\\0.5(400)(v_i^2-10^2)+(400)(9.8)(13.5)=500\\v_i=12.7\frac{m}{s}"